calculate the percentage of ionization of amobarbital (pka=8) at ph 2.0, 5.5, 8.0. What trend is seen?
Answer #1

Consider amobarbital as HA (weak acid) , Ka = 10-8

HA H+ + A-

Ka = [H+]×c/(c(1-)) = [H+] ×/(1-)

/(1-) = Ka/[H+]

At pH = 2.0 ,

/(1-) = 10-8/10-2 = 10-6

= 10-6/(1+10-6) = 9.9999×10-7

Percentage of ionization of amobarbital = 100 % = 9.9999×10-5

= 1.0×10-4 % (Answer)

At pH = 5.5 ,

/(1-) = 10-8/10-5.5 = 10-2.5 = 3.16×10-3

= 3.16×10-3/(1+3.16×10-3) = 3.15×10-3

Percentage ionization of amobarbital = 3.15×10-3 × 100%

= 0.315 % (Answer)

At pH = 8.0

/(1-) = 10-8/10-8 = 1

= 0.5

Percentage of ionization of amobarbital = 0.5×100%

= 50 % (Answer)

As the pH of solution is increased , the percentage of weak acid (amobarbital) increases because as pH increases solution becomes more basic and moves the equilibrium towards right .